A retail store had sales of $44,750 in April and $56,100 in May. The store employs eight full-time workers who work a 40-hour week. In April the store also had six part-time workers at 10 hours per week, and in May the store had nine part-timers at 17 hours per week (assume four weeks in each month).

Using sales dollars as the measure of output, what is the percentage change in productivity (dollars output per labor hour) from April to May? (Round your intermediate calculations and final answer to 2 decimal places.)
Answer #1

the average worker performance in April= 160 hours+60= 220 hours

the average worker performance= 44750/220= $203.4

the average worker performance in May= 56100/(160+153)= 56100/313= $179.23

performane in April is higher than performance in may